# Junchen Fu1 Xuri Ge2 Corresponding author. Kaiwen Zheng1 Ioannis Arapakis3 Xin Xin2 Joemon M. Jose1

### v2

@@ 5 Prompt-based Popularity Enhancement (RQ2)
(Answer to RQ2) We demonstrate the overall effectiveness of the proposed prompt enhancement, confirming its efficacy.
Video Generators Model Comparison Concrete Dataset (%) Abstract Dataset (%)
BP WR PE WR BP WR PE WR
@@
6 Benchmarking Different Models (RQ3)
@@
In this study, we employed five state-of-the-art large language models (LLMs): Llama-3.3-70B Touvron et al. (2023), Qwen-2.5-72B Yang et al. (2024a), and ChatGPT-4o Hurst et al. (2024). Additionally, we utilized the API versions of DeepSeek-V3 Liu et al. (2024a) and DeepSeek-R1 Guo et al. (2025) for evaluation. For video generation, we incorporated three widely recognized open-source models: CogVideoX-5B Yang et al. (2024b), LTX-Video HaCohen et al. (2024), and HunyuanVideo Kong et al. (2024). Furthermore, for the open-source LLMs, we adopted the instruction-tuned versions.
Table 6 demonstrates that DeepSeek models (V3 and R1) dominate most comparisons, achieving the highest number of wins. Notably, DeepSeek-R1 reaches a peak BP WR of 56% and PE WR of 66%. Meanwhile, Llama-3.3-70B stands out in PE WR for the Abstract dataset, reaching up to 56%, indicating strong adaptation to enhanced prompts in abstract reasoning. In contrast, ChatGPT-4o and Qwen-2.5-72B generally perform the worst, with Qwen-2.5-72B registering the lowest PE WR (34%) against DeepSeek-R1. These results highlight DeepSeek’s overall superiority, Llama-3.3-70B’s specific strength in abstract PE tasks, and the weaker performance of ChatGPT-4o and Qwen-2.5-72B.
Table 7 indicates that CogVideoX-5b generally underperforms compared to the other two models. In contrast, both LTX-Video and HunyuanVideo achieve win rates exceeding 50% in many pairwise comparisons against CogVideoX-5b on both the Concrete and Abstract datasets, demonstrating their relatively stronger performance. Notably, both models attain a maximum PE WR of 62% against CogVideoX-5b on the Concrete dataset. The marginal performance difference between LTX-Video and HunyuanVideo suggests that they handle diverse prompts at a comparably high level.
(Answer to RQ3) We evaluated five cutting-edge LLMs and three state-of-the-art video-generation models using the proposed pipeline. The results indicate that DeepSeek-V3 and DeepSeek-R1 rank as the top LLMs overall, consistently outperforming or tying other models in both Concrete and Abstract scenarios. For video generation, LTX-Video and HunyuanVideo deliver robust outputs with only minor gaps in win rates across different prompt types, underscoring their competitive reliability in diverse content creation tasks.
7 Related Work
Video Generation: Recent advancements in video generation have been truly breathtaking, with models such as Sora [^5] https://sora.com/ and Kling [^6] https://kling.kuaishou.com/en. Building upon the success of diffusion models in image synthesis, several studies have extended these techniques to video generation. For instance, the Make-A-Video model Singer et al. (2022) demonstrates how text-conditioned diffusion models can produce high-quality videos by jointly modeling spatial and temporal dimensions. Other approaches, such as VideoCrafter1 Chen et al. (2023a) have developed open diffusion models capable of generating high-quality videos from textual inputs, achieving resolutions up to 1024×576 pixels. Additionally, models like Control-A-Video Chen et al. (2023b) introduce controllable text-to-video generation by incorporating motion priors and content priors, enabling fine-grained control over the generated content. CogVideo Hong et al. (2022) and its successor CogVideoX Yang et al. (2024b) employ a large-scale pre-trained transformer architecture with a multi-frame-rate hierarchical training strategy, effectively aligning text and video representations for higher-quality generation. Similarly, HunyuanVideo Kong et al. (2024) introduces additional refinements, though its technical details remain limited. Further improving controllability, LTX HaCohen et al. (2024) explores novel approaches to enhance video generation fidelity.
Large Language Models for Video Generation: The emergence of large language models (LLMs) has spurred interest in their application to video generation Zhao et al. (2023); Hong et al. (2023); Maaz et al. (2023). VideoPoet Kondratyuk et al. (2023), for example, employs a decoder-only transformer architecture that processes multimodal inputs—including text, images, and videos—to generate coherent video content. Similarly, GPT4Video Wang et al. (2024) presents a unified multimodal framework that empowers LLMs with the capability of both video understanding and generation, integrating instruction-following approaches with stable diffusion models to handle various video generation scenarios.
Limitations of Existing Literature: Despite advancements in video generation and understanding, the role of large language models (LLMs) in popular video creation remains largely unexplored. Existing research primarily addresses video understanding (e.g., captioning, summarization) and video generation (e.g., script-to-video synthesis), without examining how LLMs can be effectively leveraged to enhance content popularity—a crucial factor in micro-video platforms. This gap underscores the lack of empirical studies investigating the intersection of LLMs and diffusion-based video generation in the context of audience engagement and content virality. To the best of our knowledge, we are the pioneer empirical study in this domain. This work systematically explores this underexamined area, providing foundational insights to guide future research. We aim to foster further advancements and inspire subsequent studies in AI-driven video content generation, particularly within the micro-video industry.
@@ 8 Conclusion
This empirical study demonstrates that the proposed LLMPopcorn for micro-video generation assisted by state-of-the-art LLMs performs slightly better than ordinary human-created videos, evaluated with an offline popularity predictor. These results highlight the potential of this approach. To further enhance the LLMPopcorn pipeline, we introduced prompt enhancement (PE), which exhibited strong adaptability across multiple LLMs and video-generation models. Among the tested models, DeepSeek-V3 and DeepSeek-R1 consistently performed best as LLMs, while LTX-Video and HunyuanVideo produced reliable outputs across diverse prompts.
In future work, we plan to explore reinforcement learning and fine-tuning techniques to better align prompts with contextual cues, ultimately improving video quality and engagement. This study pioneers popularity-driven micro-video generation, establishing a strong foundation for future research and innovation.
9 Limitations
Despite the promising results, this study has several limitations, which also suggest potential directions for future research in this area:
  1. 1.
    While some videos generated by LLMPopcorn using state-of-the-art LLMs exceed the median popularity of human-created content, their overall absolute scores remain modest. Nevertheless, LLMPopcorn is still able to match the average performance of human-generated videos, highlighting its potential for further refinement in aligning content with audience preferences.
  2. 2.
    We did not conduct real-world user studies to evaluate engagement with generated videos. Although the offline popularity predictor offers insights, integrating actual audience interactions (e.g., views, likes) would better assess content effectiveness.
  3. 3.
    Current state-of-the-art video generation models, despite advancements, struggle to accurately interpret complex prompts, often losing critical details and producing oversimplified or incomplete outputs.
  4. 4.
    This study does not fine-tune the LLM or video generation models, which could further enhance video quality and engagement. Incorporating popularity-based signals into video model adaptation may improve content relevance and audience resonance, potentially increasing the chances of generating viral videos.
  5. 5.
    While AI-generated videos offer significant advantages, they also present risks of misuse, such as spam content flooding social media platforms and the exploitation of recommender systems. However, most video platforms impose restrictions on account registration and daily upload limits to mitigate such risks.
We recognize that this study has several limitations, largely due to the limited existing literature in this field. However, as an initial exploration, it provides a foundation for future research. More importantly, these limitations represent opportunities for further advancements, requiring a collective effort from the broader research community. Addressing these challenges will play a crucial role in driving progress in AI-assisted popular micro-video creation.
10 Ethical Considerations
This research follows ethical guidelines. The Microlens dataset is publicly available, and all user prompts were generated by ChatGPT-4o, without human involvement or personal data. Evaluation was conducted using pre-trained models via open-source implementations or APIs, without human annotation. We strive to ensure transparency and compliance with ethical standards.
Acknowledgements
We sincerely appreciate Dr. Alexandros Karatzoglou for his invaluable guidance, which has played a crucial role in shaping the direction of this research. His insightful advice and expertise have provided significant support throughout the development of this work.
Appendix A Appendix
A.1 Implementation details
In this paper, we utilized five state-of-the-art LLMs, including two open-source models, Llama-3.3-70B and Qwen-2.5-72B, and three API-based models, DeepSeek-V3, DeepSeek-R1, and ChatGPT-4o. For efficiency, we apply 4-bit quantization to the open-source LLMs. Additionally, we incorporated three widely-used open-source video generation models: CogVideoX-5B, LTX-Video, and HunyuanVideo. All model weights were obtained from the HuggingFace platform [^7] https://huggingface.co/ to ensure consistency and reproducibility. Notably, all open-source LLMs used in this study were instruction-tuned versions. We ensured reproducibility by fixing the random seed for opensourced models, allowing for consistent results. We adopt Faiss [^8] https://github.com/facebookresearch/faiss as library for matching and indexing.
For the prompt enhancement related to RQ2, we search the retrieval-augmented generation (RAG) size from [10, 50, 100]. In terms of the embedding model, we adopt "all-MiniLM-L12-v2" [^9] https://huggingface.co/sentence-transformers/all-MiniLM-L12-v2. The video generation prompt database we adopted is from Fu et al. (2024b), which contains existing video titles and video captions paired with their popularity. All experiments are conducted on one H100 80G GPU.
A.2 License Statement
All the open-source models used in this study adhere to their respective licenses and usage terms. Specifically, Llama-3.3-70B follows Meta’s custom research license, which allows research and commercial use under specified conditions. Qwen-2.5-72B, developed by Alibaba, is released under the Tongyi Qianwen Open License, which permits research use while imposing restrictions on certain commercial applications. Similarly, the open-source video generation models, CogVideoX-5B, LTX-Video, and HunyuanVideo, are available under their respective open-access agreements, ensuring transparency and reproducibility in academic research.
For API-based models, including DeepSeek-V3, DeepSeek-R1, and ChatGPT-4o, we accessed them through their official APIs in compliance with the respective providers’ terms of use. DeepSeek models are governed by DeepSeek’s API usage policy, which allows research use while enforcing rate limits and responsible AI usage guidelines. This paper was written with ChatGPT’s assistance in compliance with OpenAI’s terms of service. The "all-MiniLM-L12-v2" embedding model is publicly available on HuggingFace under the Apache 2.0 License, allowing unrestricted research and commercial use.
By strictly following these licensing terms, we ensure that our research maintains compliance with ethical AI deployment standards, respects intellectual property rights, and upholds reproducibility for future studies.
RAG Num Concrete Abstract
LLMPopcorn 0.52 0.95
+ PE-RAG Num = 10 0.92 0.94
+ PE-RAG Num = 50 0.72 0.87
+ PE-RAG Num = 100 0.84 0.99
Table 8: Median predicted popularity for ablation study.
Figure 4: Video examples from LLMPopcorn for a concrete user query.
Figure 5: Video examples from LLMPopcorn for an abstract user query.
Figure 6: Video examples from LLMPopcorn with the Prompt Enhancement for an concrete user query.
Figure 7: Video examples from LLMPopcorn with the Prompt Enhancement for an abstract user query.
A.3 Ablation Study for the Prompt Enhancement
In this section, we analyze the impact of RAG numbers on prompt enhancement. Using DeepSeek-V3 with LTX as the default LLM and video generator, Table 8 shows that the Concrete dataset benefits more from RAG, while the Abstract dataset requires a larger scale for improvement.
For the Abstract dataset, LLMPopcorn already performs well (0.95), with only PE-RAG Num = 100 providing a notable boost (0.99). In contrast, the Concrete dataset starts lower (0.52) but becomes more effective with RAG, achieving 0.92 with PE-RAG Num = 10. This suggests that RAG may be more impactful for Concrete tasks, while Abstract tasks improve gradually with larger enhancements.
Figure 8: An overview of the Generated Video Titles and Video Prompt by LLMPopcorn.
A.4 Video and Prompt Examples
To evaluate and compare different video generation models, we apply two prompts across three models, as depicted in Figures 4, 5, 6, and 7. In these figures, the first row corresponds to videos generated by CogVideoX-5b, the second row by HunyuanVideo, and the third row by LTX-Video. The first prompt is abstract: "Create a video about the history and evolution of a popular street food." The second is a concrete prompt: "Create a video explaining the world-building in Avatar (2009)."
A.5 Efficiency
We present Table 9 and Table 10, which report the efficiency of three key processes: retrieval for PE, LLM inference for LLMPopcorn and PE, video generation, and evaluation. Efficiency is measured in terms of VRAM usage and inference time. All evaluations are conducted on a single H100 GPU.
VRAM Time
Retrieval Process 896 12 s/query
Video Generation
CogVideoX 28,788 170 s/video
Hunyuan 12,016 54 s/video
LTX 27,100 20 s/video
Evaluation Process 3,000 4 s/video
Table 9: Efficiency of the pipeline procedures. We denote Time as the inference time (measured in seconds, denoted as $s$), and VRAM is measured in megabytes (MB).
LLM Pipeline VRAM (MB) Inference Time (s/query)
Qwen-2.5-72B LLMPopcorn 44,246 8
+PE-RAG Num = 10 47,468 28
+PE-RAG Num = 50 53,504 62
+PE-RAG Num = 100 67,192 90
LLaMA-3.3-70B LLMPopcorn 42,240 9
+PE-RAG Num = 10 43,948 48
+PE-RAG Num = 50 49,662 75
+PE-RAG Num = 100 56,842 86
Table 10: Inference efficiency test results for LLMPopcorn on Qwen-2.5-72B and LLaMA-3.3-70B
Figure 9: An overview of the basic prompt template for LLMPopcorn.
Figure 10: An overview of the Prompt Enhancement template.

### v3

@@ 4 Popular Micro-videos Generation (RQ1)
We evaluate three LLMs—Llama-3.3-70B, ChatGPT-4o, and DeepSeek-V3—with CogVideoX as the default video generation model, using concrete and abstract prompt datasets from Section 3. We compare LLMPopcorn-generated video popularity to human-created videos from the Microlens dataset [20] by matching prompts to semantically similar video titles using an embedding model to reduce thematic bias. Table 2 shows DeepSeek-V3-generated videos achieve median popularity scores of 0.56 (concrete) and 0.73 (abstract), rivaling Microlens’ 0.44 (concrete) and 0.71 (abstract). However, absolute scores remain low, likely due to limitations in open-source video models [18]. Figure 3 reveals DeepSeek-V3’s smoother popularity distribution for abstract prompts, with more videos in mid-to-high ranges, while Llama-3.3-70B and ChatGPT-4o show concentrated low scores with rare outliers. For concrete prompts, model differences are minimal. Table 1 indicates DeepSeek-V3 leads in abstract dataset categories like Anime, Film, and Delicacy, with less dominance in Daily Sharing and Game. (Answer to RQ1) LLMs enhance micro-video creation by structuring descriptions, with DeepSeek-V3 achieving popularity comparable to or surpassing human-created videos, highlighting AI’s potential for engaging short video generation.
5 Popularity Enhancement (RQ2)
Video Generators Comparison Concrete Abstract
  BP PE  BP PE
@@ 5 Popularity Enhancement (RQ2)
Table 3 compares Prompt Enhancement (PE) to the basic LLMPopcorn prompt using win rates from pairwise comparisons across models. LLama-3.3-70B achieves 56% (concrete) and 52% (abstract) win rates, outperforming LLMPopcorn. DeepSeek-V3 scores 55% (concrete) and 49% (abstract), while ChatGPT-4o maintains 53% (abstract), showing PE’s creative compatibility. A user study with DeepSeek-V3 evaluated PE against LLMPopcorn. Participants blindly compared five video pairs per prompt type, selecting the more popular video. Table 4 shows PE achieved over 60% win rates for both prompt types. (Answer to RQ2) PE significantly enhances prompt effectiveness, validated by offline metrics and user studies.
6 Benchmarking Models (RQ3)
@@
We evaluated five large language models (LLMs): Llama-3.3-70B [22], Qwen-2.5-72B [26], ChatGPT-4o [13], DeepSeek-V3, and DeepSeek-R1 [9], using their API versions. For video generation, we tested CogVideoX-5B [27], LTX-Video [11], and HunyuanVideo [16], with instruction-tuned versions for open-source LLMs.
Table 5 shows DeepSeek-V3 and DeepSeek-R1 leading, with DeepSeek-R1 achieving a peak BP WR of 56% and PE WR of 66%. Llama-3.3-70B performs strongly in PE WR for abstract tasks (56%), while Qwen-2.5-72B lags with a PE WR of 34% against DeepSeek-R1. Table 6 indicates LTX-Video and HunyuanVideo outperform CogVideoX-5B, both exceeding 50% win rates in pairwise comparisons, with a peak PE WR of 62% on the Concrete dataset. (Answer to RQ3) DeepSeek-V3 and DeepSeek-R1 excel among LLMs, consistently leading in Concrete and Abstract scenarios. LTX-Video and HunyuanVideo show comparable, reliable performance in video generation across diverse prompts.
Fig. 4: Video examples from LLMPopcorn, shown from left to right: abstract user queries (1–3) followed by concrete user queries (4–6).
Fig. 5: Video examples from LLMPopcorn with Prompt Enhancement, shown from left to right: abstract user queries (1–3) and concrete user queries (4–6).
7 Visualization
To compare different video generation models, we present visualizations generated using two prompts across three models, as shown in Figures 4 and 5. Each figure includes three rows: videos in the first row are generated by CogVideoX-5b, those in the second row by HunyuanVideo, and those in the third row by LTX-Video. The first prompt is abstract: “Create a video about the history and evolution of a popular street food.” The second prompt is a concrete prompt: “Create a video explaining the world-building in Avatar (2009).” These examples help illustrate how each model handles abstract versus concrete prompts. In all figures, the first three columns represent abstract data, while the last three represent concrete data. From the two figures, Prompt Enhancement may boost popularity by enriching queries to produce dynamic, narrative-driven content. This is shown in its ability to transform abstract prompts into a more diverse range of imaginative scenes (whereas the baseline results are more uniform), while also depicting concrete subjects through engaging actions, like a chef’s hands at work, instead of as static objects.
@@ 8 Conclusion
This empirical study demonstrates that the proposed LLMPopcorn for micro-video generation assisted by state-of-the-art LLMs performs slightly better than ordinary human-created videos, evaluated with an offline popularity predictor. These results highlight the potential of this approach. To further enhance the LLMPopcorn pipeline, we introduced prompt enhancement (PE), which exhibited strong adaptability across multiple LLMs and video-generation models. Among the tested models, DeepSeek-V3 and DeepSeek-R1 consistently performed best as LLMs, while LTX-Video and HunyuanVideo produced reliable outputs across diverse prompts. In future work, we plan to explore reinforcement learning and fine-tuning techniques to better align prompts with contextual cues, ultimately improving video quality and engagement. This study pioneers popularity-driven micro-video generation, establishing a strong foundation for future research and innovation.